Blog from January, 2017

v04 Install

Tower and soil array installed, unable to power up and test due to lack of power drop for site. Post-install an EC100 box grounding issue was raised. Additional grounding connecting CSAT 3A sensor head, EC100, and tower must be installed. Soil serial cable is too short, will either reinstall soil within 5m to the tower or determine if there is enough 15m serial cables to install one at this site.

To Do:

  • EC100 / CSAT 3A ground
  • soil serial to dsm cable
  • run power and test sensors when drop is installed
v05 Install

Installed sensors and powered up v05 without incident. Although it's relay (rnw06) is not up yet, it was able to establish a network connection with another router and transmit. 

Left to do at site:

  •  hose clamp grounding wire to tower

 

Daily status 31 Jan

A lot of progress today:

  • The tower tension-checkers determined that all 10 and 20m towers should be safe to climb.
  • AC power was turned on to 2 towers: rsw01 and rsw02.  rsw01's WiFi was reprogrammed, though the access point for both towers is not yet up (soon?)
  • 2 stations (v04 and v05) were instrumented by us and also are powered on(?).
  • Another 8 stations were prepped at the the base of the mast, to be ready for the INEGI crew to climb tomorrow and move the instrumentation vertically.  There are a few details still to take care of at these masts, including preparing solar panels at 2 sites and installing soil sensors at 2 sites, but we expect that this will be quick.
  • 2 more stations were prepped at the ops center.
  • The power company was supposed to have finished at rsw06 as well, so that we can bring this up as a repeater site for the western slope (including ISS).
  • DTU fixed DSM power issues at tnw07, tnw10, (and tse07?)
  • By my count, this leaves only 4 masts that are constructed but not yet prepped at the tower.  We still think that these 4 can be finished Thurs & Fri.

 

 

 

tse07 CSAT back up

Power at tse07 is now at 12.13V, so I think it is high enough to power the lower CSAT back up.

 

rsw05 staged at tower
Ted and Gary staged sensors for rsw05.
  • Missing nut for samortecnica boom cross brace.
  • Need to plug A/D port with puttty.
  • The setra needs to be moved, it is not clear enough of the cross-brace and not quite at 2m.  Move it up to the right height and off the lacing.
The seismometer needs to plug into power aux port on DSM.  The picklist says seismometer plugs into a/d channel 2, but that was an earlier offer that will not be used.  The seismometer has its own logger.
Serial numbers:
Metek: 0102122233

 

tnw09 staged at tower

Gary and Ted setup sensors on the tower in preparation for INEGI climbers the next day.

Inventory at site:
  • Ec150 1392
  • Csat3a 2035
  • Nr01 2536
  • EC100 box 1693
  • RMYoung ARL 98c0056-038
  • Ubiquiti
  • DSM tnw09
  • 2 x 15m bulgin cables
  • Soil sensors had already been installed, so those serial numbers were not recorded by us.

 

tnw03 staged at tower
Inventory at tower:
  • 5m cable, Metek s/n 2007 09007/01
  • 10m cable, Metek s/n 2007 08017/01
  • Ubiquiti
  • Mote ID5

Ethernet access port will need to be plugged with putty.

Meteks aimed with arrow towards tower along boom, lettering facing away.

 

Daily status 30 Jan

Some movement today...

INEGI will send two crews here on Wed to get as much done as possible.  The prediction of rain and other INEGI commitments have prevented them from committing to more time, though Thurs might be possible.  In the meantime, our tensiometers have arrived and we have learned that we can climb if our staff determine that a particular tower is safe.  We are optimistic that this criterion can be met for the majority of the towers.  Tomorrow, we will deliver as much equipment as possible to the site and prepare booms to be raised in order to maximize tower climbing time.

Other tidbits:

  • Gordon has fixed the biggest bug in cockpit, so we now have real-time plots
  • A few more towers have been prepped in the ops center
  • Electricians have been working on the sw ridge, so we connected DC power to 2 towers (though the main line is not yet live)
  • There have been some power issues in the ops center itself, including the breaker for the ballroom tripping.  We yet don't know the source.

 

tse13 TRH problems

on tse13, the 60 and 80m TRHs weren't showing up.  80m just needed a ^R to restart the fan.  (I note that RH=100%.)  60m seems totally dead, even after a power cycle. (sad)

 

Found that the "differential protection" circuit for the ops center was tripped when I got here at 1830.  From what I see in the data files, this might have happened about 1630, with ustar losing data (but of course, the DSMs stilll haven't been rsync'd).

 

tse13t mote turned off

I realized that, if the cable to the mote is disconnected, it wasn't a good idea for it to have live power.  Thus, I did a "pio 7 0" this morning (we had light rain then).  This should be installed in a dsm_up script so that it is repeated if power is cycled.

  • Check tse13t power mote, and if necessary reconnect, and turn the port power back once connection is confirmed.
Daily status 28 Jan

A few things done today.  Now all sensors are reporting from the 10 towers that are operating.

  • We've organized our teams and instrumentation to be able to start immediately on Monday.
  • DTU again came to the rescue and installed the missing SHT sensors in our TRH housings on tse13.  In the process, we've lost the voltage-monitoring mote at tse13t, but this isn't a big deal.  I'm guessing that the Binder cable was knocked off.
  • We installed DC power drops at 3 valley sites.
  • 3 more towers are prepped.
  • tse09's 24V power connection came loose this morning and was serviced.
  • I played around a bit more with webplots.
  • Kurt and I did feasibilities for 3 future projects...

Consistent with most of VERTEX and Perdigao, tomorrow (Sunday) is a down day.

 

Daily status 27 Jan

Some progress being made...

Our tensiometers were held up in customs.  This now should be solved, but they won't arrive until Monday.  In the meantime, we have learned that the document we had on tower tensions a) isn't for any of the towers being used here and b) has a specification for the maximum tension for the guy wires, but not the working tension.  We will ask if there are other documents and may be able to go back to the engineering firm that prepared this document to determine a working tension.  However, we now are in a situation where we do not have a set of specifications to compare any of our measured tensions to.  In this case, we do not think that we will (ever) be allowed to climb these towers.

Thus, we have asked INEGI if they can supply climbers for all of the towers, hopefully starting next week.  Obviously, we would provide all ground support.  We will meet tomorrow to see if we can organize into 3 crews, each with one or two INEGI climbers.  A secondary issue is that we would need INEGI climbers for all maintenance visits and tear-down as well.

Now for the good news: DTU has enabled AC power now for 4 of the tall masts.  After some work on the network, software updates for the DSMs that we hung in December, and enabling of DTU data, we now have 10 sites reporting.  There are two more 60m masts that require minor AC power fixes, but this needs to be done by the electricians.

Other items:

  • All towers that have been built are now prepped
  • 3 of the NCAR team, Cornell, and 1 of the DTU team have now headed home

 

Daily status 26 Jan

Another slow day, especially with rain that has come and is forecast to hang around for a while.  Most of the crew worked a half day.  Tensiometers still hadn't arrived by 6pm tonight, so we're anxious about tomorrow and the weekend.  I hope to update this shortly with news that they are here.  Otherwise, we'll have nothing to do until Monday...

However, due to DTU's efforts, we now have power to tse13.  Since this is a repeater, this allowed connection to several other towers that have been powered in the last few days.  Thus, we now have data from 6 towers streaming in.  At least one more tower is powered up and just needs a few WiFi settings to be adjusted.

In other news:

  • Finished prepping the valley sites (even the 3 towers that are not yet erected).  Now only have 1 more tower erect that hasn't been prepped.
  • Got the rudiments of wwwplots running.  This is being served up to the world at: http://ops-perdigao.dyndns.org/isf/projects/Perdigao/isfs/qcdata/  Scripts still need to be started to update this page automatically.
  • Apparently, Gordon has been updating cockpit, though we haven't started running it yet (due to no streams of data until late today).
  • The last (of 3) 4-component radiometer installations and its soil sensors are done.  We still have about 10 soils to install, but many are waiting for Samortecnica to finish work at these sites (since there has been a significant amount of surface disturbance as the towers are built).
  • Ted and Santiago were to have finished configuration of the project's WiFi gear.
  • Cornell leaves tomorrow, having installed their seismometers and 3 of 4 lidars.
  • DTU will have help from the municipality to install their 2 or 3 heavy lidars tomorrow.

 

RMY odd

S/N 01256 (from OU) seems not to like being touched.  In the lab, it works fine sitting down, but when touched flags nearly every sample as bad.  We should put this mounted low, just in case this problem occurs in the field.